OpenStudy (anonymous):

let R be the set of real numbers and d :RxR be a metric on R defined by d(x,y)=|x-y|. obtain d(x,2x)

OpenStudy (zzr0ck3r):

\(d(x,2x)=|x-2x|=|-x|=|x|\)

OpenStudy (anonymous):

thank you sir

OpenStudy (anonymous):

can i then say that |x|= norm of x which is (x*x)^1/2 ?

OpenStudy (anonymous):

i have another question sir'

OpenStudy (anonymous):

\[\let x=R^2 and d \infty : XxX---> R on X define by d \infty(X,Y)=MAX lei \le2|\Xi,Yi| for all X=(x1-x2), Y=(y1-y2). obbtain d \infty(X,Y) if X=(-1,2) and Y=(3,-4)\]

OpenStudy (zzr0ck3r):

I can't read that. You can always make a metric a norm.

OpenStudy (anonymous):

OK. thank u sir